In​ March, 2006 the labor department of the country of Jobland determined that the size of the labor force was

102102

million out of which

9696

million were employed. In May​ 2006, the labor department determined that

1.51.5

million people became​ "discouraged workers" in April.

Assuming that no new jobs were created between March and​ April, calculate the unemployment rate for March and April.

Jobland unemployment rate in March​ =

nothing

percent.

​(Enter

your response rounded to one decimal​ place.)
Answer #1

Answer:

March 2006

Labor force = 102 million

Employed = 96 million

April 2006

Discouraged workers = 1.5 million

No new jobs were created between March and April.

In March 2006, Unemployed = 102 million - 96 million = 6 million

Unemployment rate in March 2006 = (102-96)/102 = 0.0588 or 5.9%

Discouraged workers in April 2006 = 1 million - This number is not to be counted as unemployed nor included in labor force as they are considered non-participant in the labor market

Unemployment rate in April 2006 = 102 million - 96 million – 1.5 million = 4.5 million

Labor force in April 2006 = 102 million – 1.5 million = 100.5 million

Hence, unemployment rate in April 2006 = 4.5/100.5 = 4.5%